# Self-regularized entropy: How much do quantum black holes violate the no-hair theorem?

### v1

@@ VI Summary and conclusions
We have shown that a single geometric deformation; the small, positive static quadrupole encoded by the $q$-metric removes the ultraviolet divergence of the entropy of Hawking radiation that plagues the brick wall treatment of Schwarzschild. In the extreme near-horizon region the angular metric acquires anisotropic scaling, $g_{\theta\theta}\propto\rho^{2q}$ and $g_{\phi\phi}\propto\rho^{-2q}$, which modifies the centrifugal barrier and converts the Schwarzschild logarithm into a convergent power. Counting quasinormal modes in the WKB approximation and performing both the radial overtone and angular sums in closed form, we obtained a compact expression for the canonical entropy that tends to the usual logarithmic divergence in the limit $q$ approaching zero.
Combining that microscopic result with the macroscopic requirement that the entropy trapped within the angular momentum barrier does not exceed the Bekenstein-Hawking value sets the lower bound of $|q|\simeq 0.18$-$0.19$ for astrophysical black holes. In other words, the distortion that cures the brick wall is not arbitrarily small: thermodynamic self-consistency fixes its natural scale.
Although our derivation used the static $q$-metric, the same parameter controls clean diagnostics in a stationary extension. In particular, multipole combinations that vanish for Kerr acquire sensitivity to $q$ (thus violating the no-hair theorem), while near-horizon lensing observables (photon ring shape and centroid) are perturbed at the percent-to-ten percent level for $|q|\,\sim\,0.2$. These signatures are natural targets for the ngEHT and gravitational wave spectroscopy of EMRIs with LISA. In short, observationally testable quadrupole distortion thus links ultraviolet finiteness, gravitational wave echo phenomenology, and forthcoming observations of supermassive black holes, offering a concrete bridge between horizon-scale quantum physics and astronomical measurements. On the theory front, extending the analysis to spinning backgrounds and/or additional field content, as well as quantifying backreaction and waveform systematics for LISA, are natural next steps.

### v2

@@ VI Summary and conclusions
We have shown that a single geometric ingredient (a sufficiently small but nonzero static quadrupolar deformation of the Schwarzschild exterior encoded by the $q$-metric) removes the brick-wall ultraviolet divergence in the canonical entropy of a probe scalar field at any finite asymptotic temperature. In the near-horizon limit the transverse geometry becomes anisotropically scaled, which modifies the effective centrifugal barrier and converts the Schwarzschild logarithmic mode-counting divergence into a convergent power law. Using WKB counting of trapped near-horizon cavity modes (together with the large-$\ell$ angular density approximation), we obtained a closed-form expression for the canonical entropy, which reduces to the standard brick-wall divergence as $q\to 0$.
Matching this canonical entropy $S(q)$ to the macroscopic Bekenstein-Hawking entropy scale $S_{\rm BH}$ (with the thermodynamic inputs specified earlier) yields an entropy-motivated deformation scale. For astrophysical entropies $S_{\rm BH}\sim 10^{78}\text{-}10^{97}$ this gives $|q|\simeq 0.18$-$0.19$, within the small-$|q|$ regime in which the near-horizon expansion is controlled. In other words, the distortion that cures the brick wall is not arbitrarily small: thermodynamic self-consistency fixes its natural scale. In this benchmark, the deformation needed to cure the brick-wall divergence is therefore modest but not parametrically tiny.
Although our entropy calculation is performed in the static geometry, the stationary extension of the $q$-metric provides a convenient phenomenological map from $q$ to measurable multipole deviations: Kerr-null combinations such as $Q_{1}$ and $Q_{2}$ acquire $O(|q|)$ values. Taking the entropy-motivated scale $|q|\sim 0.2$ as a fiducial target, this corresponds to percent-to-tens-of-percent departures from the Kerr multipole relations for moderate spins, and motivates sensitivity studies with ngEHT strong-field lensing observables, EMRI multipole measurements by LISA, and spin-induced quadrupole in 3G ground-based gravitational wave observatories.
On the theory front, natural next steps include extending the mode count to spinning black holes, treating the angular spectrum beyond the large-$\ell$ approximation, incorporating additional field content, backreaction/renormalization, performing end-to-end ray-tracing and EMRI waveform studies that propagate these deviations into realistic observational inference pipelines.